The one-sample t statistic for testing

H0: μ = 20

Ha: μ < 20

based on n = 7 observations has the value

t = −1.89.

(a)

What are the degrees of freedom for this statistic?

(b)

Between what two values does the P-value of the test fall? (You may use Table D.)

A) 0.005 < P < 0.010

B) .01 < P < 0.02    

C) 0.02 < P < 0.025

D) 0.025 < P < 0.05

E) 0.05 < P < 0.1

(c)

If you have software available, find the exact P-value. (Round your answer to four decimal places.)
Answer #1

A) Degrees of freedom = (n-1) = 7-1 = 6

One sample student 't' statistic follows 't' distribution with (n-1) degrees of freedom

B) (E)

0.05 < P < 0.1

C) P-Value = 0.0538 (one-tail) at '6' degrees of freedom